# Simultaneous Khintchine theorem on manifolds in positive characteristics: convergence case

### v1

@@ 4.3. Analysis on Special Part
Note that in the context of Theorem 4.5, up to a constant $\widetilde{C}$, we have
$r=t,s^{\prime}=\frac{s-t}{2},t_{1}=\cdots=t_{n}=s.$
Hence,
$\begin{split}s^{\prime}+\sum_{i}t_{i}-r-\max_{i}t_{i}=(n-1)s-t+\frac{s-t}{2}=\frac{2n-1}{2}s-\frac{3}{2}t<0,\end{split}$
whenever $t$ is large enough. By Theorem 4.5 and Lemma 4.6, there exist $C>0$ and $\alpha>0$ such that
@@
$\begin{split}\mathcal{L}_{d}\left(\mathfrak{M}(s,t)\cap B\right)\leq&C\max\left\{q^{-t},q^{\frac{(2n-1)s-3t}{2(n+1)}}\right\}^{\alpha}\mathcal{L}_{d}(B).\end{split}$
@@
By combining Theorem 2.3, Lemma 4.3, and Lemma 4.7 we obtain Proposition 1.
5. Proof of Khintchine’s theorem: converegence case
Before proceeding to the proof of Theorem 1.2, we prove the following fact, which will be useful later in the proof of Theorem 1.2.
Lemma 5.1.
Let $\psi:(0,\infty)\to(0,1)$ be a non-increasing function. Then
$\sum_{Q\in\mathbb{F}_{q}[T]\setminus\{0\}}\psi(|Q|)^{n}<\infty\quad\Longleftrightarrow\quad\sum_{t=0}^{\infty}q^{t}\psi(q^{t})^{n}<\infty.$
Proof.
We know that every nonzero polynomial $Q\in\mathbb{F}_{q}[T]$ can be uniquely written as
$Q=a\cdot Q_{0},$
where $a\in\mathbb{F}_{q}^{\times}$ is a nonzero scalar and $Q_{0}$ is a monic polynomial. Furthermore, there are exactly $q^{t}$ monic polynomials of degree $t$, and each such $Q$ satisfies $|Q|=q^{t}$. Therefore,
$\sum_{Q\in\mathbb{F}_{q}[T]\setminus\{0\}}\psi(|Q|)^{n}=(q-1)\sum_{\begin{subarray}{c}Q\in\mathbb{F}_{q}[T]\setminus\{0\}\\ Q\text{monic}\end{subarray}}\psi(|Q|)^{n}=(q-1)\sum_{t=0}^{\infty}\sum_{\begin{subarray}{c}Q\text{ monic}\\ \deg Q=t\end{subarray}}\psi(q^{t})^{n}=(q-1)\sum_{t=0}^{\infty}q^{t}\psi(q^{t})^{n}.$
We are now in a position to prove Theorem 1.2.
Proof of Theorem 1.2.
Without loss of generality, it suffices to prove that for any $\mathbf{x}_{0}\in U$ and any small ball $B_{0}$ around $\mathbf{x}_{0}$
$\mathcal{L}_{d}\left(\left\{\mathbf{x}\in B_{0}:\mathbf{f}(\mathbf{x})\in\mathcal{S}_{n}^{\bm{\Theta}}(\psi)\right\}\right)=0,$
whenever $\psi$ is monotonic and
$\sum_{Q\in\mathbb{F}_{q}[T]\setminus\{0\}}\psi(|Q|)^{n}<\infty.$
@@ 5. Proof of Khintchine’s theorem: converegence case
We divide the estimate into two parts: the special part, i.e., $\mathfrak{M}(-\log_{q}\psi(q^{t}),t)$, and the generic part, which comprises the set $B_{0}\setminus\mathfrak{M}(-\log_{q}\psi(q^{t}),t)$. Now we make the following observation. If $\mathbf{f}(\mathbf{x})\in\mathcal{S}_{n}^{\bm{\Theta}}(\psi)$, then there exist infinitely many $t\in\mathbb{N}$ such that
@@
$\left\|\mathbf{f}(\mathbf{x})-\frac{{\mathbf{P}}+\bm{\Theta}}{Q}\right\|\leq\frac{\psi(q^{t})}{q^{t}}\quad\text{with}\,\,(\mathbf{P},Q)\in\mathcal{R}^{n+1}\,\,\text{and}\,\,|Q|=q^{t}.$
Hence, for every $T\geq 1$, we have
@@
$\begin{split}\left\{\mathbf{x}\in B_{0}:\mathbf{f}(\mathbf{x})\in\mathcal{S}_{n}^{\bm{\Theta}}(\psi)\right\}\subseteq\bigcup_{t\geq T}\underbrace{\mathfrak{M}(-\log_{q}\psi(q^{t}),t)\cap B_{0})}_{A_{t}^{\bm{\Theta}}}\bigcup\\ \bigcup_{t\geq T}\underbrace{\bigcup_{\tiny(\mathbf{P},Q)\in\mathcal{R}_{\bm{\Theta}}(B_{0}\setminus\mathfrak{M}(-\log_{q}\psi(q^{t}),t))}\left\{\mathbf{x}\in B_{0}:\left\|\mathbf{x}-\frac{\pi(\mathbf{P}+\bm{\Theta})}{Q}\right\|\leq\frac{\psi(q^{t})}{q^{t}}\right\}}_{B_{t}^{\bm{\Theta}}},\end{split}$ (5.1)
@@
where $\pi:\mathcal{K}^{n}\to\mathcal{K}^{d},$ is the projection into the first $d$ coordinates. To bound the measure of $B_{t}^{\bm{\Theta}}$, we invoke Proposition 2 to obtain
$\begin{split}\mathcal{L}_{d}(B_{t}^{\bm{\Theta}})\leq q^{t(d+1)}\psi(q^{t})^{m}\frac{\psi(q^{t})^{d}}{q^{td}}=q^{t}\psi(q^{t})^{n}.\end{split}$ (5.2)
By Lemma 5.1, the series $\sum_{t\geq T}\psi(q^{t})^{n}q^{t}$ converges. Therefore, by (5.2), $\lim_{T\rightarrow\infty}\mathcal{L}_{d}\left(\bigcup_{t\geq T}B_{t}^{\Theta}\right)=0$. For the other part, that is,
$A_{t}^{\bm{\Theta}}:=\mathfrak{M}(-\log_{q}\psi(q^{t}),t)\cap B_{0},$
we apply Lemma 4.7 to estimate the measure of the set. Since $\lim_{|Q|\rightarrow\infty}\psi(|Q|)=0$, we may assume that $\psi(q^{t})\geq q^{-\frac{t}{2n-1}}$ for all $t$ sufficiently large. Otherwise, one can replace $\psi(q^{t})$ by $\max\{\psi(q^{t}),q^{-\frac{t}{2(n-1)}}\}$. Thus, $\log_{q}\psi(q^{t})\leq\frac{t}{2n-1}$ for every $t$ large enough. By Lemma 4.7, there exists $\alpha>0$ such that either
@@
$\begin{split}\mathcal{L}_{d}\left(\mathfrak{M}(-\log_{q}\psi(q^{t}),t)\cap B_{0})\right)&\leq Cq^{\alpha\left(-\frac{2n-1}{2(n+1)}\log_{q}\psi(q^{t})-\frac{3}{2(n+1)}t\right)}\mathcal{L}_{d}(B_{0})\\ &\leq Cq^{-\frac{\alpha}{(n+1)}t}\mathcal{L}_{d}(B_{0}),\end{split}$ (5.3)
or,
$\mathcal{L}_{d}\left(\mathfrak{M}(-\log_{q}\psi(q^{t}),t)\cap B_{0})\right)\leq Cq^{-\alpha t}\mathcal{L}_{d}(B_{0}).$ (5.4)
Thus, by (5.2), (5.3) and (5.4), we have
$\sum_{t\geq T}\mathcal{L}_{d}(B_{t}^{\bm{\Theta}})+\sum_{t\geq T}\mathcal{L}_{d}(\mathfrak{M}(-\log_{q}\psi(q^{t}),t)\cap B_{0})<\infty,$
@@
so that by the Borel-Cantelli lemma, $\mathcal{L}_{d}(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi)))=0$. ∎
5.1. Proof of Hausdorff Dimension
Proof of Theorem 1.3.
@@ 5.1. Proof of Hausdorff Dimension
Take $s=-\log_{q}\psi(q^{t})>0$. By Theorem Proposition 1 and 2.3, one can cover $\mathfrak{M}(s,t)$ with disjoint balls of radius $q^{-\frac{s+t}{2}}$, such that the number of balls is $\ll q^{\frac{d}{2}(s+t)}\mathcal{L}_{d}(\mathfrak{M}(s,t))$. Thus by Proposition 1, there exists $T_{0}\geq 1$ such that for every $t\geq T_{0}$, we have
$\mathcal{H}^{\sigma}(A_{t}^{\bm{\Theta}})\ll q^{-\frac{s+t}{2}(\sigma-d)}\mathcal{L}_{d}(\mathfrak{M}(s,t))\ll q^{-\frac{s+t}{2}(\sigma-d)}.$ (5.5)
Hence, by plugging in $s=-\log_{q}\psi(q^{t})$,
[…]
For the set $B_{t}^{\bm{\Theta}}$, we employ the counting estimate (4.2) along with a covering by balls of radius $r=q^{-t}\psi(q^{t})$. Hence there exists $T_{1}\geq 1$ such that
$\sum_{t\geq T_{1}}\mathcal{H}^{\sigma}(B_{t}^{\bm{\Theta}})$ $\ll\sum_{t\geq T_{1}}q^{(d+1)t}q^{-ms}r^{\sigma}$
$=\sum_{t\geq T_{1}}{\psi(q^{t})}^{\sigma+m}q^{(d+1)t-\sigma t}$
[…]
Therefore, combining both for $t\geq T:=\max\{T_{0},T_{1}\}$, we obtain
@@
$\sum_{t\geq T}\mathcal{H}^{\sigma}(A_{t}^{\bm{\Theta}})+\sum_{t\geq T}\mathcal{H}^{\sigma}(B_{t}^{\bm{\Theta}})<\infty.$ (5.7)
As a consequence, $\mathcal{H}^{\sigma}(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi)))=0$. ∎
@@ 6. Appendix - A Result in Geometry of Numbers for Discrete Subgroups
We consider an extension of the function field $\mathcal{K}$ obtained by adjoining the power $T^{\frac{1}{\ell}}$. As mentioned earlier, we denote this extension by $\mathcal{K}_{\ell}$. Similarly, denote the lattice $\mathbb{F}_{q}[T^{1/\ell}]\subseteq\mathcal{K}_{\ell}$ by $\widetilde{\mathcal{R}}:=\mathbb{F}_{q}[T^{\frac{1}{\ell}}]$. We now prove the following proposition, which establishes a relationship between the first minimum of the discrete subgroup $gu\mathcal{R}^{m+n}\subset\mathcal{K}_{\ell}^{m+n}$ and that of the lattice $gu\widetilde{\mathcal{R}}^{m+n}\subset\mathcal{K}_{\ell}^{m+n}$, where $g$ is some diagonal matrix in $\operatorname{SL}(m+n,\mathcal{K}_{\ell})$ and $u$ is some unipotent matrix in $\operatorname{SL}(m+n,\mathcal{K})$.
Proposition 3.
Let $g\in\operatorname{SL}\left(m+n,\mathcal{K}_{\ell}\right)$ be a diagonal matrix, and let $\bm{\alpha}\in M_{m\times n}(\mathcal{K})$ and
@@
Let $(\mathbf{P},\mathbf{Q})=(P_{1},\dots,P_{m},Q_{1},\dots,Q_{n})^{T}\in\widetilde{\mathcal{R}}^{m+n}$ be a vector. Explicitly write $P_{i}=\sum_{k=0}^{l-1}T^{\frac{k}{\ell}}P_{i,k}$, where $P_{i,k}\in\mathcal{R}$, and $Q_{j}=\sum_{k=0}^{\ell-1}T^{\frac{k}{\ell}}Q_{j,k}$, where $Q_{j,k}\in\mathcal{R}$. Since every $\alpha\in\mathcal{K}_{\ell}$ can be written as $\alpha=T^{\deg(\alpha)}\cdot\frac{\alpha}{T^{\deg(\alpha)}}$, where $\deg(\alpha)\in\frac{1}{\ell}\mathbb{Z}$ and $\left|\frac{\alpha}{T^{\deg(\alpha)}}\right|=1$, there exist $r_{1},\dots,r_{m+n}\in\mathbb{Z}$ and $\mathbf{o}=\operatorname{diag}\{o_{1},\dots,o_{m+n}\}$, with $|o_{j}|=1$ for every $j=1,\dots,m+n$, such that
@@
$\left\|gu\begin{pmatrix}\mathbf{P}\\ \mathbf{Q}\end{pmatrix}\right\|=\max_{k=0,\dots,\ell-1}q^{\frac{k}{\ell}}\left\|\begin{pmatrix}P_{1,k}+\sum_{j=1}^{n}\alpha_{1,j}Q_{j,k}\\ \vdots\\ P_{m,k}+\sum_{j=1}^{n}\alpha_{m,j}Q_{j,\ell}\\ Q_{1,k}\\ \vdots\\ Q_{n,k}\end{pmatrix}\right\|$. (6.2)
@@
Therefore, $\lambda_{1}\left(gu\widetilde{\mathcal{R}}^{m+n}\right)\geq\lambda_{1}\left(gu\mathcal{R}^{m+n}\right)$. On the other hand, $gu\mathcal{R}^{m+n}\subseteq gu\widetilde{\mathcal{R}}^{m+n}$, and thus, $\lambda_{1}\left(gu\widetilde{\mathcal{R}}^{m+n}\right)=\lambda_{1}\left(gu\mathcal{R}^{m+n}\right).$
@@
By Proposition 3, the set introduced in equation (4.1) can equivalently be written as
@@
Proposition 3 gives rise to several questions about the comparison between successive minima of discrete subgroups in extension fields and lattices in extension fields. Hence, we conclude this section with the following questions.

### v2

@@ 4.3. Analysis on the Special Part
Note that, using Lemma 4.6, in the context of Theorem 4.5, up to a constant $\widetilde{C}$, we have
$r=t,s^{\prime}=\lceil{(s-t)/2}\rceil,t_{1}=\cdots=t_{n}=s.$
For $t\gg 1$, clearly $s^{\prime}+\sum_{i}t_{i}-r-\max_{i}t_{i}<(n-1)s-t+\frac{s-t}{2}+1<1+ns-\frac{3t}{2}\leq 1-\left(\kappa-\frac{3}{2}\right)t<0$, holds for all $s\in\mathbb{N}$ such that $s\leq\frac{\kappa t}{n}$.
By Theorem 4.5 and Lemma 4.6, for every $\mathbf{x}_{0}\in U$ there exists a ball $B$ centered at $\mathbf{x}_{0}$ and constants $\hat{C},\alpha>0$ such that
@@
$\begin{split}\mathcal{L}_{d}\left(\mathfrak{M}(s,t)\cap B\right)\leq\hat{C}\max\left\{q^{-t},q^{\left(\frac{2n-1}{2(n+1)}s-\frac{3t}{2(n+1)}+1\right)}\right\}^{\alpha}\mathcal{L}_{d}(B)&=\hat{C}q^{\alpha\left(\frac{2n-1}{2(n+1)}s-\frac{3t}{2(n+1)}+1\right)}\mathcal{L}_{d}(B)\\ &\leq\left(\hat{C}q^{\alpha}\right)q^{\frac{\alpha}{n+1}\left(ns-\frac{3t}{2}\right)}\mathcal{L}_{d}(B).\end{split}$
@@
By combining Theorem 2.3, Lemma 4.3, and Lemma 4.7 we obtain Proposition 1.1. We now proceed to prove the main theorems.
5. Proofs of Theorem 1.2 and Theorem 1.4
5.1. Proof of Theorem 1.2
To begin with, we remind the reader that the approximating function $\psi$ has the decay rate specified in (1.4). It is enough to show that for any $\mathbf{x}_{0}\in U$ and any small ball $B_{0}\ni\mathbf{x}_{0}$, one has
$\mathcal{L}_{d}\left(\left\{\mathbf{x}\in B_{0}:\mathbf{f}(\mathbf{x})\in\mathcal{S}_{n}^{\bm{\Theta}}(\psi)\right\}\right)=0.$
@@
We divide the manifold into special and generic parts, i.e., $\mathfrak{M}(-\log_{q}\psi(q^{t}),t)$ and $B_{0}\setminus\mathfrak{M}(-\log_{q}\psi(q^{t}),t)$. To this end, we first observe that if $\mathbf{f}(\mathbf{x})\in\mathcal{S}_{n}^{\bm{\Theta}}(\psi)$, then there exist infinitely many $t\in\mathbb{N}$ such that
@@
$\left\|\mathbf{f}(\mathbf{x})-\frac{{\mathbf{P}}+\bm{\Theta}}{Q}\right\|\leq\frac{\psi(q^{t})}{q^{t}}\quad\text{with}\,\,(\mathbf{P},Q)\in\mathcal{R}^{n}\times(\mathcal{R}\setminus\{0\})\,\,\text{and}\,\,|Q|=q^{t}.$
Hence, for every $N\in\mathbb{N}$, we have
@@
$\begin{split}\left\{\mathbf{x}\in B_{0}:\mathbf{f}(\mathbf{x})\in\mathcal{S}_{n}^{\bm{\Theta}}(\psi)\right\}\subseteqq\bigcup_{t\geq N}\underbrace{\mathfrak{M}(-\log_{q}\psi(q^{t}),t)\cap B_{0})}_{A_{t}^{\bm{\Theta}}}\bigcup\\ \bigcup_{t\geq N}\underbrace{\bigcup_{\tiny(\mathbf{P},Q)\in\mathcal{R}_{\bm{\Theta}}(B_{0}\setminus\mathfrak{M}(-\log_{q}\psi(q^{t}),t))}\left\{\mathbf{x}\in B_{0}:\left\|\mathbf{x}-\frac{\pi(\mathbf{P}+\bm{\Theta})}{Q}\right\|\leq\frac{\psi(q^{t})}{q^{t}}\right\}}_{B_{t}^{\bm{\Theta}}},\end{split}$ (5.1)
@@
where $\pi:\mathcal{K}^{n}\to\mathcal{K}^{d},$ is the projection into the first $d$ coordinates. We invoke Proposition 4.1 to obtain
$\begin{split}\mathcal{L}_{d}(B_{t}^{\bm{\Theta}})\leq q^{t(d+1)}\psi(q^{t})^{m}\frac{\psi(q^{t})^{d}}{q^{td}}=q^{t}\psi(q^{t})^{n},\penalty 10000\ \forall t\gg 1.\end{split}$ (5.2)
Lemma 4.7 will be applied to estimate the measure of $A_{t}^{\bm{\Theta}}$. Since (1.4) is equivalent to
$\log_{q}\psi(q^{t})\leq\frac{\kappa t}{n},\text{ for all }t\in\mathbb{N},$
Lemma 4.7 yields an $\alpha>0$ such that, for all sufficiently large $t\in\mathbb{N}$,
@@
$\begin{split}\mathcal{L}_{d}\left(A_{t}^{\bm{\Theta}}\right)\leq Cq^{\frac{\alpha}{n+1}\left(-n\log_{q}\psi(q^{t})-\frac{3t}{2}\right)}\mathcal{L}_{d}(B)\leq Cq^{\frac{\alpha}{(n+1)}\left(\kappa-\frac{3}{2}\right)t}\mathcal{L}_{d}(B_{0}).\end{split}$ (5.3)
From (5.2), (5.3) and ($\mathfrak{R}$.1), we conclude that
$\sum_{t\geq N}\mathcal{L}_{d}(B_{t}^{\bm{\Theta}})+\sum_{t\geq N}\mathcal{L}_{d}\left(A_{t}^{\bm{\Theta}}\right)<\infty,\text{ when $N$ is large enough}.$
@@
$\mathcal{L}_{d}(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi)))=0$ is now immediate from this in view of the Borel-Cantelli lemma.
5.2. Proof of Theorem 1.4
@@
Take $s=-\log_{q}\psi(q^{t})>0$. By Proposition 1.1 and Theorem 2.3, one can cover $A_{t}^{\bm{\Theta}}=\mathfrak{M}(s,t)\cap B_{0}$ with disjoint balls of radius $q^{-\frac{s+t}{2}}$. When $t\in\mathbb{N}$ is large enough, one can cover it with at most $q^{\frac{d}{2}(s+t)}\mathcal{L}_{d}(A_{t}^{\bm{\Theta}})$ such balls. Thus, by Proposition 1.1, there exists $N_{1}\in\mathbb{N}$ such that for every $t\geq N_{1}$, we have
$\mathcal{H}^{\sigma}(A_{t}^{\bm{\Theta}})\ll q^{-\frac{s+t}{2}(\sigma-d)}\mathcal{L}_{d}(A_{t}^{\bm{\Theta}})\ll q^{-\frac{s+t}{2}(\sigma-d)}q^{{\frac{\alpha}{n+1}}(ns-\frac{3t}{2})}.$ (5.4)
Hence, by plugging in $s=-\log_{q}\psi(q^{t})$, one obtains from (1.5) and ($\mathfrak{R}$.1) that
@@
$\sum_{t\geq N_{1}}\mathcal{H}^{\sigma}(A_{t}^{\bm{\Theta}})\ll\sum_{t\geq N_{1}}\left(\frac{\psi(q^{t})}{q^{t}}\right)^{\frac{\sigma-d}{2}}\left(\psi(q^{t})^{n}q^{\frac{3t}{2}}\right)^{-\frac{\alpha}{n+1}}<\infty.$ (5.5)
For the set $B_{t}^{\bm{\Theta}}$, we employ the counting estimate (4.2) along with a covering by balls of radius $r=q^{-t}\psi(q^{t})$. Hence, there exists $N_{2}\in\mathbb{N}$ such that
$\sum_{t\geq N_{2}}\mathcal{H}^{\sigma}(B_{t}^{\bm{\Theta}})\ll\sum_{t\geq N_{2}}q^{(d+1)t}q^{-ms}r^{\sigma}=\sum_{t\geq N_{2}}\left({\frac{\psi(q^{t})}{q^{t}}}\right)^{\sigma+m}q^{(n+1)t}<\infty,$ (5.6)
owing to (1.5) and ($\mathfrak{R}$.1). It follows at once from (5.1), (5.5) and (5.6) that $\mathcal{H}^{\sigma}(\mathbf{f}^{-1}(\mathcal{S}_{n}^{\bm{\Theta}}(\psi)))=0$, due to the Borel-Cantelli lemma.
@@ 6. Appendix - A Result in Geometry of Numbers for Discrete Subgroups
We consider an extension of the function field $\mathcal{K}$ obtained by adjoining the power $T^{\frac{1}{\ell}}$. As mentioned earlier, we denote this extension by $\mathcal{K}_{\ell}$. Similarly, we denote the lattice $\mathbb{F}_{q}[T^{1/\ell}]\subseteqq\mathcal{K}_{\ell}$ as $\widetilde{\mathcal{R}}{\,\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\,}\mathbb{F}_{q}[T^{\frac{1}{\ell}}]$. We now prove the following proposition, which establishes a relationship between the first minimum of the discrete subgroup $gu\mathcal{R}^{m+n}\subseteqq\mathcal{K}_{\ell}^{m+n}$ and that of the lattice $gu\widetilde{\mathcal{R}}^{m+n}\subseteqq\mathcal{K}_{\ell}^{m+n}$, where $g$ is some diagonal matrix in $\operatorname{SL}(m+n,\mathcal{K}_{\ell})$ and $u$ is some unipotent matrix in $\operatorname{SL}(m+n,\mathcal{K})$.
Proposition 6.1.
Let $g\in\operatorname{SL}\left(m+n,\mathcal{K}_{\ell}\right)$ be a diagonal matrix, $\bm{\alpha}\in M_{m\times n}(\mathcal{K})$ and let
@@
Let $(\mathbf{P},\mathbf{Q})=(P_{1},\dots,P_{m},Q_{1},\dots,Q_{n})^{T}\in\widetilde{\mathcal{R}}^{m+n}$ be a vector. Explicitly write $P_{i}=\sum_{k=0}^{\ell-1}T^{\frac{k}{\ell}}P_{i,k}$, where $P_{i,k}\in\mathcal{R}$, and $Q_{j}=\sum_{k=0}^{\ell-1}T^{\frac{k}{\ell}}Q_{j,k}$, where $Q_{j,k}\in\mathcal{R}$. Since every $\alpha\in\mathcal{K}_{\ell}$ can be written as $\alpha=T^{\deg(\alpha)}\cdot\frac{\alpha}{T^{\deg(\alpha)}}$, where $\deg(\alpha)\in\frac{1}{\ell}\mathbb{Z}$ and $\left|\frac{\alpha}{T^{\deg(\alpha)}}\right|=1$, there exist $r_{1},\dots,r_{m+n}\in\mathbb{Z}$ and $\mathbf{o}=\operatorname{diag}\{o_{1},\dots,o_{m+n}\}$, with $|o_{j}|=1$ for every $j=1,\dots,m+n$, such that
@@
$\left\|gu\begin{pmatrix}\mathbf{P}\\ \mathbf{Q}\end{pmatrix}\right\|=\max_{k=0,\dots,\ell-1}q^{\frac{k}{\ell}}\left\|\begin{pmatrix}P_{1,k}+\sum_{j=1}^{n}\alpha_{1,j}Q_{j,k}\\ \vdots\\ P_{m,k}+\sum_{j=1}^{n}\alpha_{m,j}Q_{j,k}\\ Q_{1,k}\\ \vdots\\ Q_{n,k}\end{pmatrix}\right\|$. (6.2)
@@
Therefore, $\lambda_{1}\left(gu\widetilde{\mathcal{R}}^{m+n}\right)\geq\lambda_{1}\left(gu\mathcal{R}^{m+n}\right)$. On the other hand, $gu\mathcal{R}^{m+n}\subseteqq gu\widetilde{\mathcal{R}}^{m+n}$, and thus, $\lambda_{1}\left(gu\widetilde{\mathcal{R}}^{m+n}\right)=\lambda_{1}\left(gu\mathcal{R}^{m+n}\right).$
@@
By Proposition 6.1, the set introduced in equation (4.1) can equivalently be written as
@@
Proposition 6.1 gives rise to several questions about the comparison between successive minima of discrete subgroups in extension fields and lattices in extension fields. Hence, we conclude this section with the following questions.